OpenStudy (anonymous):

How do I show this result: There are n 1's.. 1111111..(n times) Then if n = p * m where p and m are integers => 11....1 (n 1's) = (11...1)*(100..0100...01......1), where the right factor in the right side is composed of 1, then p-1 0's, then 1, then p-1 0's, etc., up to the last 1 which is the m-th one.

OpenStudy (abb0t):



OpenStudy (anonymous):

@amistre64 , @Yahoo!

OpenStudy (abb0t):

A line over the top of the repeated number(s)

OpenStudy (anonymous):

I dont understand??

OpenStudy (anonymous):

Could you refer to the q again?

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

this is not always true. eg 111 = 37x3

OpenStudy (experimentx):

maybe 11111... = 99999../9 = (10^(n+1)-1)/9 ... If this n+1 is not prime, then (10-1)(10+1)(100+1) ... /9 something like that ... sorry, extremely busy due to exams!!

OpenStudy (anonymous):

1111111111111

OpenStudy (anonymous):

@Ineedanswersbadfail ?? @experimentX It meant that 111, has 3 ones. So 3=3x1. Therefore it can only be expressed as 111*1

OpenStudy (experimentx):

That's a trivial case when you the (number of digits +1) is prime ... I think we could use above idea. Suppose (the number of digits +1)is not prime. \[ 1111...111 \text{ (m times)}(m+1 = p q) \\ = \frac{10^{m+1}-1}{9}=\frac{10^{pq}-1}{9} = \frac{10^p - 1}{9}(1 + 10^q + 100^q+ ... + 10^{(p-1)q}) \\ = 111...111 \times 10010...01001\]